OpenStudy (anonymous):

true or false. the y intercept of a straight line with equation Ax + by +c = 0 is C/B (B to the power of 1 0)

OpenStudy (raden):

false that should -c/b

OpenStudy (anonymous):

yes it should thak you. why is it false

OpenStudy (anonymous):

thank

OpenStudy (raden):

to get y-intercept, subtitute x = 0 to line ax + by + c = 0 a(0) + by + c = 0 0 + by + c = 0 by + c = 0 by = 0 - c by = -c y = -c/b you'rer welcome

OpenStudy (anonymous):

thank you again so much for your time